# Scalp Diagnostic System With Label-Free Segmentation and Training-Free Image Translation

### v2

@@ 3.3 Scalp Condition Classification
4 Experiments
To showcase the effectiveness of our pipeline, we conducted a series of quantitative and qualitative experiments for each component of the pipeline.
4.1 Dataset
We trained and evaluated our model using the scalp image dataset from AI-Hub [1], previously mentioned in Section 1 and depicted in Figure 1. The dataset comprises $95{,}910$ images with a resolution of $640\times 480$ pixels from $20{,}000$ patients. It is divided into $72{,}342$ images for training and $23{,}568$ images for testing, with a subset of $21{,}703$ images from the training set designated for validation. Dermatologists classified these scalp images into one or more of the following conditions: dandruff, excess sebum, and erythema. Each condition was further assessed for severity, categorized as mild, moderate, or severe, with an additional category, good, indicating no presence of the condition. Furthermore, a subset of $23{,}803$ images, divided into $12{,}902$ for training, $5{,}613$ for validation, and $5{,}288$ for testing, includes specific labels for alopecia severity, following the same four-level categorization.
@@ 4.2 Hair Segmentation
Table 1: Performance of Hair segmentation on manually annotated test set.
Approach Pixel-F1 Jaccard Dice
Shih et al. [44] $0.7063$ $0.3475$ $0.5119$
Yue et al. [52] $0.7943$ $0.4933$ $0.6540$
Kim et al. [23] $0.8150$ $0.5614$ $0.7078$
SAM [26] $0.5030$ $0.3609$ $0.5024$
Ours $\hat{M}$ $0.8534$ $0.6039$ $0.7481$
$M_{\text{AP}}$ $0.8358$ $0.5951$ $0.7429$
$M$ 0.8680 0.6494 0.7859
To evaluate the performance of the hair segmentation methods, we manually annotated hair regions in $150$ images from the test set. For baseline methods, we selected previous studies related to scalp image analysis  [44, 52, 23] and SAM [26]. Additionally, we performed an ablation study comparing our final segmentation result ($M$) with the intermediate results ($\hat{M}$ and $M_{\text{AP}}$).
Quantitative results. Table 1 reveals that our methods surpass the performance of existing hair segmentation techniques. In particular, the approach of combining the advantages of the two masks, $\hat{M}$ and $M_{\text{AP}}$ using the AND operator in $M$ showed the best performance. These results show the limitations of traditional computer vision techniques used in previous studies for image segmentation, revealing a lack of understanding in capturing the intricate patterns of hair and the scalp. Additionally, we observed that SAM, when used without specific guidance, was less effective for automatic image segmentation.
@@
4.3 Scalp Disease and Severity Diagnosis
Figure 6: Image translation results with different generative models. The goal is to maintain the content of the source image while translating the scalp condition to the target image.
Figure 7: Image translation results using various mask guidance. Note that our approach is guided by $1-M$.
Table 2: Performance of scalp condition classification with various augmentation methods, denoted after “+” symbol, on the test set. The second column displays the overall macro-F1 score, while the third to last column shows the F1 scores for each severity level of three diseases.
@@ 4.3 Scalp Disease and Severity Diagnosis
DenseNet [21] $0.5819$ $0.7962$ $0.5140$ $0.5916$ $0.6138$ 0.5541 $0.6013$ $0.6412$ $0.0000$ 0.7758 $0.7287$ $0.5652$ $0.6010$
+ Gaussian Noise $0.5667$ $0.7804$ $0.4969$ $0.5662$ $0.5971$ $0.4712$ $0.5809$ $0.5946$ $0.0000$ $0.7509$ $0.7119$ $0.6142$ $0.6352$
+ AugMix [18] $0.5249$ $0.7886$ $0.5009$ $0.5890$ $0.0000$ $0.5042$ $0.5882$ $0.6341$ $0.0000$ $0.7430$ $0.7179$ $0.5956$ $0.5845$
+ DiffuseIT [27] $0.6079$ $0.8090$ $0.4824$ $0.6041$ $0.6500$ $0.5357$ $0.6126$ $0.6245$ $0.2024$ $0.7744$ $0.7395$ $0.6209$ $0.6388$
+ AGG [28] $0.6103$ $0.8106$ $0.4803$ $0.5908$ $0.6539$ $0.5175$ $0.5977$ $0.6119$ $0.2997$ $0.7713$ 0.7402 0.6291 $0.6206$
+ Ours 0.6358 0.8196 0.5407 0.6253 0.6652 $0.5362$ 0.6174 0.6414 0.4296 $0.7578$ $0.7339$ $0.6214$ 0.6412
EfficientFormerV2 [30] $0.5692$ $0.7954$ $0.4172$ $0.5983$ $0.6279$ $0.5257$ $0.5648$ $0.6283$ $0.0000$ $0.7722$ $0.7090$ $0.6227$ $0.5685$
+ Gaussian Noise $0.5618$ $0.7799$ $0.4774$ $0.5658$ $0.6326$ $0.4601$ $0.5850$ $0.5501$ $0.0000$ $0.7418$ $0.7141$ $0.5980$ $0.6372$
+ AugMix [18] $0.5770$ $0.7885$ $0.4940$ $0.5916$ $0.6347$ $0.5189$ $0.5927$ $0.6228$ $0.0000$ $0.7460$ $0.7238$ $0.6199$ $0.5904$
+ DiffuseIT [27] $0.5955$ $0.7981$ $0.4410$ $0.5977$ $0.6316$ $0.5263$ $0.5953$ $0.6057$ $0.2362$ $0.7658$ $0.7149$ $0.6121$ $0.6214$
+ AGG [28] $0.6098$ $0.8012$ $0.5087$ $0.6037$ $0.6262$ $0.5112$ $0.5830$ $0.6084$ $0.2999$ 0.7874 $0.7364$ $0.6239$ $0.6275$
+ Ours 0.6346 0.8066 0.5290 0.6190 0.6689 0.5345 0.6129 0.6324 0.4060 $0.7811$ 0.7382 0.6394 0.6476
Synthetic image generation. For the evaluation of DiffuseIT-M, we compared our model against DiffuseIT [27] and AGG [28] as baseline for the state-of-the-art image-to-image translation model. We have selected to employ the FID [19] and LPIPS [55] scores for fidelity evaluation. Comparative experiments were conducted using images from our augmentation dataset, with DiffuseIT and AGG serving as baseline models. Table 3 reveals that DiffuseIT-M outperforms other models in both metrics, indicating superior image fidelity. This high-quality image generation is attributed to our model’s effective implementation of mask guidance. As shown in Figure 6, we can observe that both DiffuseIT and AGG models fail to preserve the hair content information from the source image. Furthermore, these models tended to compromise overall information and were unable to transfer the semantic information. However, our model successfully preserved hair content information and transferred the semantic information.
Table 3: Quantitative analysis of image-to-image translation performance.
Model FID ($\downarrow$) LPIPS ($\downarrow$)
DiffuseIT [27] $138.42$ $0.4631$
AGG [28] $141.70$ $0.4915$
Ours 74.84 0.3528
Effect of mask guidance. We conducted experiments to examine the impact of mask guidance on hair information preservation during image translation. As illustrated in Figure 7, our method, guided by the mask $1-M$, effectively retains hair features while successfully transferring the semantic attributes of the target image onto the scalp. In contrast, using the reverse mask, $M$, leads to only minor alterations in scalp color from the target image, with a notable transfer of hair semantic information from the target. When no mask ($\mathbf{0}$) is applied, the translation results in minimal color change, failing to transfer conditions like dandruff from the target image. Conversely, with a full mask ($\mathbf{1}$), both hair and scalp features are subjected to changes. This differentiation in results highlights the importance of mask guidance in preserving specific image features, demonstrating the versatility of our approach in handling different translation objectives.
Scalp condition classification. To demonstrate the effectiveness of our augmentation method using generated images, we employed two different models as the classification backbone: DenseNet [21] as a CNN and EfficientFormerV2 [30] as a Transformer. We conducted a comparative analysis with other augmentation methods, including the addition of Gaussian noise, AugMix [18], DiffuseIT [27], and AGG [28]. The results are summarized in Table 2.
@@
Our approach, specifically employing DiffuseIT-M, achieved the highest performance in both models. Notably, classifying the severe sebum class proved to be especially challenging when using non-generative augmentation methods. This difficulty arises primarily due to the extreme scarcity of samples for this class. The augmentation of the training dataset with generative models led to enhanced performance compared to the baseline, thanks to the effective style and content guidance from these models. Our model, in particular, exhibited superior accuracy compared to DiffuseIT and AGG, which struggled to preserve the essential information of the hair effectively. This underscores the significance of incorporating both the scalp style details and the hair content information in scalp disease classification.
4.4 Alopecia Severity Diagnosis
Table 4: Performance of alopecia severity classification.
Method Accuracy Weighted F1
Image-Based DenseNet [51] $0.6686$ $0.6283$
EfficientFormerV2 [12] $0.6728$ $0.6351$
Feature-Based Kim et al. [23] $0.5994$ $0.6062$
Ours 0.7460 0.6873
Figure 8: Distribution of features used for alopecia prediction. This figure illustrates the distribution of $\mathcal{T}$, representing hair thickness values, and $|B|$, indicating the count of bounding boxes in each image.
To diagnose the severity of alopecia, we utilized hair-related features such as hair thickness and count, as described in Section 3.2. These features were then employed to make predictions using a GradientBoosting model [15].
Our method was benchmarked against both image-based and traditional feature-based diagnostic approaches. For image-based severity prediction, we used DenseNet and EfficientFormerV2 models as baselines, which directly analyze the images to predict alopecia severity. Conversely, for feature-based prediction, we employed a method from [23], which involves generating masks through Otsu-based thresholding [34] and extracting features based on average hair thickness.
Since the data on alopecia severity is imbalanced, as shown in Figure 1, we selected to use both accuracy and weighted F1-score as evaluation metrics.
Alopecia severity classification. Our approach consistently outperforms other methods, including those based on image analysis, as shown in Table 4. This highlights the significance of hair thickness and counts as key diagnostic features for alopecia and validates the accuracy of our method in predicting these attributes precisely.
Feature distribution. We analyzed the distribution of input features across different alopecia severity levels. As illustrated in Figure 8, box plots for each feature show a consistent decrease from the good to severe categories. This trend indicates that our balanced sample set allowed each feature to contribute effectively to the severity prediction, enhancing the model’s accuracy.
Table 5: Comparative analysis of different prompting methods with SAM. The symbol “–” indicates scenarios where no prompt was used.
Positive Negative Pixel-F1 Jaccard Dice
$\hat{M}$ $1-\hat{M}$ $0.3786$ $0.1199$ $0.1891$
^$\hat{\mathcal{H}}_{\text{skel}}$ $1-\hat{M}$ $0.8210$ $0.5707$ $0.7215$
$\hat{C}$ $0.4966$ $0.2754$ $0.4148$
$\hat{C}$ $1-\hat{M}$ 0.8358 0.5951 0.7429
5 Ablation Studies
5.1 Automatic Prompting for SAM
We conducted an ablation study to evaluate our automatic prompting method with SAM. The study comprised four different approaches: 1) positive point prompts from $\hat{M}$ and negative point prompts from $1-\hat{M}$, 2) positive point prompts from ^$\hat{\mathcal{H}}_{\text{skel}}$ and negative point prompts from $1-\hat{M}$, 3) positive point prompts from $\hat{\mathcal{C}}$ without negative prompts, and 4) positive prompt points from $\hat{\mathcal{C}}$ and negative point prompts from $1-\hat{M}$ (Ours).
Quantitative results. Table 5 demonstrates that our method outperforms the others across all metrics. Notably, the absence of negative prompts significantly diminishes performance, underscoring their importance in the model.
Figure 9: Illustration of hair segmentation results using different prompt guides. “GT” represents segmentation guided by the ground truth mask, while $M_{\text{AP}}$ signifies our final prompting method with positive prompts from $\hat{\mathcal{C}}$ and negative prompts from $(1-\hat{M})$. “(+)(+)( + )” refers to positive prompts and “()(-)( - )” refers to negative prompts.
Qualitative results. Figure 9 showcases the efficacy of our point prompting method in hair segmentation. While using ^$\hat{\mathcal{H}}_{\text{skel}}$ for prompting improves over $\hat{M}$, it often results in unevenly distributed sampled points, leading to a noisier mask than that produced by our method.
5.2 Hair Preservation Efficacy in Generative Models
We conducted an analysis to understand how various image-to-image generative models impact hair segmentation. Specifically, we transformed 150 test images by applying a healthy target image (devoid of any scalp diseases) using three different generative models. We then applied our segmentation method to these images for a thorough quantitative evaluation.
Quantitative results. For a numerical evaluation of hair preservation, we generated images and then applied segmentation metrics like pixel F1, Jaccard index, and Dice score. Table 6 showcases the results, highlighting that DiffuseIT-M (denoted as Ours ($M$)) excels in preserving hair details, thereby affirming its efficacy in retaining essential content from the source images. Moreover, we examined the segmentation outcomes when using GT masks (denoted as Ours (GT)) alongside our model. This comparison illustrated that improving mask quality significantly enhances the preservation of hair features in the generated images, providing insights into how mask quality impacts hair content retention in generative models.
Qualitative results. We compared hair preservation in images generated by our model with those from other models. Our model used GT masks as guides during the generation process. Figure 10 shows the hair segmentation results for images generated by our model alongside the GT masks. The comparison demonstrates that our model effectively retains hair content from source images, unlike DiffuseIT and AGG, which struggled to maintain comprehensive hair details. In particular, AGG tended to preserve only dominant hair features, omitting finer details.
Table 6: Comparative analysis of hair segmentation across different generative methods.
Model Pixel-F1 Jaccard Dice
DiffuseIT $0.7106$ $0.3619$ $0.5153$
AGG $0.7286$ $0.3785$ $0.5403$
Ours ($M$) $0.7625$ $0.4485$ $0.6067$
Ours (GT) 0.7812 0.4840 0.6416
Figure 10: Comparative qualitative results of hair segmentation across different generative models.
Table 7: Ablation study of impact analysis of features in the alopecia severity prediction. $HC$ is the number of Hair estimation from Section 3, $\mathcal{T}$ is the list of hair thickness estimations and $|B|$ is the number of bounding boxes in an image. The “ $\tilde{\mathcal{T}}$ ” symbol denotes the median of $\mathcal{T}$, while the “ $\overline{\mathcal{T}}$ ” symbol signifies the mean of $\mathcal{T}$. An $\ast$” symbol signifies that outliers were removed. The order in which features are added reflects feature importance in the final model.
$HC$ $\tilde{\mathcal{T}}^{\ast}$ $\tilde{\mathcal{T}}$ $|B|$ $\overline{\mathcal{T}}$ $\overline{\mathcal{T}}^{\ast}$ Accuracy F1 (Weighted)
$0.7233$ $0.6109$
$0.7369$ $0.6650$
$0.7390$ $0.6712$
$0.7407$ $0.6761$
$0.7436$ $0.6783$
0.7460 0.6873
5.3 Feature Selection for Severity Estimation
We identified key features for alopecia severity prediction, focusing on hair thickness and count. For thickness, we analyzed the length of the list $B$ and computed the mean and median values of $\mathcal{T}$, excluding outliers identified using the interquartile range (IQR). We also considered the mean and median of the entire $\mathcal{T}$. Additionally, the number of line segments detected by the hair count estimator was used as a supplementary feature.
Feature importance. An analysis using the GradientBoosting model (Table 7) revealed that the best prediction accuracy was achieved when all features were combined. This finding underscores the significance of each extracted feature in our model.
6 Conclusion
In this work, we introduced ScalpVision, diagnostic system designed for a complete evaluation of scalp health. This system harnesses label-free learning combined with a foundational model to create precise hair segmentation masks. Additionally, ScalpVision employs diffusion-based image augmentation for accurate scalp condition diagnosis and leverages estimations of hair thickness and count to assess alopecia severity. We envision ScalpVision as a significant step towards a generalized diagnostic system for dermatological applications, particularly in simplifying the challenges associated with complex labeling in microscopic scalp imaging.
7 Limitation
There are several limitations of ScalpVision that would make for promising avenues of future work, which mainly steams from dataset constraints. The limited labels in the dataset pose challenges in accurately diagnosing scalp and alopecia conditions. For example, while follicle count is crucial for alopecia diagnosis, we used hair count as a proxy due to data limitations. However, not all visible hairs in an image necessarily originate from follicles within the same image, making this a less than ideal substitute. Additionally, we assumed that scalp images capture a consistent area. In reality, slight variations in the distance from which images are captured can lead to differences in perceived hair thickness. Addressing these discrepancies requires further research into standardizing image capture methods or developing models robust to such variations.
Acknowledge
@@ 7 Limitation
This work was partly supported by an IITP grant (No. RS-2020-II201361, RS-2022-00155966) and NRF grant (No. RS-2024-00354218) funded by the Korea government (MSIT).
Supplementary Material
This supplementary material provides additional details and extended results that complement the main paper. Included in this supplementary material are the following sections:
  • Section A delves into detailed explanations of various medical scalp diseases.
  • Section B offers a comprehensive look at our experimental setup and provides supplementary information.
  • Section C introduces pseudo training set for training the hair segmentation model.
  • Section D presents further insights into our approach to scalp disease classification.
  • Section E extends the discussion on the selection and analysis of features used in predicting the severity of alopecia.
  • Section F shows additional qualitative results of DiffuseIT-M.
Appendix A Detailed Overview of Scalp Diseases
The dataset from AI-Hub [^3] https://aihub.or.kr categorizes scalp images into three primary conditions: dandruff, excess sebum, and erythema.

### v3

@@ 3.1 Dataset
Table 1: Performance of hair segmentation on the test set.
Approach Pixel-F1 Jaccard Dice
Shih et al. [28] $0.706$ $0.348$ $0.512$
Yue et al. [32] $0.794$ $0.493$ $0.654$
Kim et al. [13] $0.815$ $0.561$ $0.708$
SAM [15] $0.503$ $0.361$ $0.502$
Ours $\hat{M}$ $0.853$ $0.604$ $0.748$
$M_{\text{AP}}$ $0.836$ $0.595$ $0.743$
$M$ 0.868 0.649 0.786
Table 2: Quantitative analysis of image-to-image translation.
Model FID ($\downarrow$) LPIPS ($\downarrow$)
DiffuseIT $138.42$ $0.463$
AGG $141.70$ $0.492$
Ours 74.84 0.353
@@ 3.2 Hair Segmentation
Because narrow hairs are difficult to segment, many existing unsupervised methods still rely on traditional computer vision for hair-specific challenges. Accordingly, we compare prior scalp segmentation approaches [28, 32, 13] as baselines and also evaluate the foundational model SAM [15]. In addition, we conduct an ablation study on our final mask $M$ and its intermediate versions $\hat{M}$ and $M_{\mathrm{AP}}$.
Quantitative results. Table 2 reveals that our methods surpass the performance of existing hair segmentation techniques. In particular, the approach of combining the advantages of the two masks, $\hat{M}$ and $M_{\text{AP}}$ using the logical AND operator in $M$ showed the best performance. These results show the limitations of traditional computer vision techniques used in previous studies for image segmentation, revealing a lack of understanding in capturing the intricate patterns of hair and the scalp. Additionally, SAM was less effective for automatic segmentation when used without specific guidance.
@@
3.3 Synthetic Image Generation
For the evaluation of DiffuseIT-M, we compared our model against DiffuseIT [16] and AGG [17] as baselines for the image-to-image translation model. This evaluation demonstrates that our model not only achieves high fidelity in image-to-image translation but also effectively preserves the desired hair details.
Figure 4: Image translation results with different generative models, where the goal is to preserve source hairlines while changing the scalp.
Figure 5: Image translation results using various mask guidance. Note that our approach is guided by 1M1-M.
Quantitative results. We have selected to employ the FID [11] and LPIPS [33] scores for fidelity evaluation using images from our augmentation dataset, with DiffuseIT and AGG serving as baseline models. Table 2 reveals that DiffuseIT-M outperforms other models in both metrics, indicating superior image fidelity. This high-quality image generation is attributed to our model’s effective implementation of mask guidance.
Qualitative results. Figure 5 shows that both DiffuseIT and AGG models fail to preserve the hair content information from the source image. Furthermore, these models tended to compromise overall information and were unable to transfer the semantic information. However, our model successfully preserved hair content information and transferred the semantic information.
Effect of mask guidance. We conducted experiments to examine the impact of mask guidance on hair information preservation during image translation. As illustrated in Figure 5, our method, guided by the mask 1M1-M, effectively retains hair features while successfully transferring the semantic attributes of the target image onto the scalp. In contrast, using the reverse mask, $M$, leads to only minor alterations in scalp color from the target image, with a notable transfer of hair semantic information from the target. When no mask ($\mathbf{0}$) is applied, the translation results in minimal color change, failing to transfer conditions like dandruff from the target. Conversely, with a full mask ($\mathbf{1}$), both hair and scalp features are subjected to changes. This differentiation in results highlights the importance of mask guidance in preserving specific image features, demonstrating the versatility of our approach in handling different translation objectives.
Table 3: Performance of scalp condition classification with various augmentation methods, denoted after “+” symbol, on the test set. The second column displays the overall macro-F1 score, while the columns from the third onward show the F1 scores for each severity level of the three diseases.
@@ 3.3 Synthetic Image Generation
DenseNet $0.582$ $0.796$ $0.514$ $0.592$ $0.614$ 0.554 $0.601$ $0.641$ $0.000$ 0.776 $0.729$ $0.565$ $0.601$
+ Gaussian Noise $0.567$ $0.780$ $0.497$ $0.566$ $0.597$ $0.471$ $0.581$ $0.595$ $0.000$ $0.751$ $0.712$ $0.614$ $0.635$
+ AugMix $0.525$ $0.789$ $0.501$ $0.589$ $0.000$ $0.504$ $0.588$ $0.634$ $0.000$ $0.743$ $0.718$ $0.596$ $0.585$
+ DiffuseIT $0.608$ $0.809$ $0.482$ $0.604$ $0.650$ $0.536$ $0.613$ $0.625$ $0.202$ $0.774$ $0.740$ $0.621$ $0.639$
+ AGG $0.610$ $0.811$ $0.480$ $0.591$ $0.654$ $0.518$ $0.598$ $0.612$ $0.300$ $0.771$ 0.740 0.629 $0.621$
+ Ours 0.636 0.820 0.541 0.625 0.665 $0.536$ 0.617 0.641 0.430 $0.758$ $0.734$ $0.621$ 0.641
EfficientFormerV2 $0.569$ $0.795$ $0.417$ $0.598$ $0.628$ $0.526$ $0.565$ $0.628$ $0.000$ $0.772$ $0.709$ $0.623$ $0.569$
+ Gaussian Noise $0.562$ $0.780$ $0.477$ $0.566$ $0.633$ $0.460$ $0.585$ $0.550$ $0.000$ $0.742$ $0.714$ $0.598$ $0.637$
+ AugMix $0.577$ $0.789$ $0.494$ $0.592$ $0.635$ $0.519$ $0.593$ $0.623$ $0.000$ $0.746$ $0.724$ $0.620$ $0.590$
+ DiffuseIT $0.596$ $0.798$ $0.441$ $0.598$ $0.632$ $0.526$ $0.595$ $0.606$ $0.236$ $0.766$ $0.715$ $0.612$ $0.621$
+ AGG $0.610$ $0.801$ $0.509$ $0.604$ $0.626$ $0.511$ $0.583$ $0.608$ $0.300$ 0.787 $0.736$ $0.624$ $0.628$
+ Ours 0.635 0.807 0.529 0.619 0.669 0.535 0.613 0.632 0.406 $0.781$ 0.738 0.639 0.648
3.4 Scalp Condition Classification
@@
To demonstrate the effectiveness of our augmentation method using generated images, we employed two different models as the classification backbone: DenseNet  [12] as a CNN and EfficientFormerV2 [18] as a Transformer. As summarized in Table 3, we compared only those augmentation methods that preserve a one-to-one correspondence between each image and its original, unambiguous set of condition labels – Gaussian noise, AugMix [10], DiffuseIT [16], and AGG [17]. Our approach, which specifically employs DiffuseIT-M, achieved the highest performance in both models. Notably, classifying the severe sebum class proved to be especially challenging when using non-generative augmentation methods. This difficulty arises primarily due to the extreme scarcity of samples for this class. The augmentation of the training dataset with generative models led to enhanced performance compared to the baseline. Our model, in particular, exhibited superior accuracy compared to DiffuseIT and AGG, which struggled to preserve the essential information of the hair effectively. This underscores the significance of incorporating both the scalp style details and the hair content information in the scalp disorder classification.
4 Conclusion and Discussion
In this work, we introduced ScalpVision, a diagnostic system designed for a complete evaluation of scalp health. Our approach combines label-free hair segmentation – based on a naive segmentation model and a foundation segmentation model (SAM) – with diffusion-based data augmentation to address data imbalance and preserve critical hair features. However, scalp disorders are affected by both the condition of the scalp and hair characteristics. Therefore, we plan to incorporate hair information to broaden our research to conditions such as alopecia, beyond the three scalp diseases. We see ScalpVision as an important step toward a more general diagnostic system for dermatological applications.
{credits}
4.0.1 Acknowledgements
@@
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2024-00354218 and RS-2024-00353125). Junhyug Noh was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2022-00155966).
Appendix 0.A Detailed Overview of Scalp Diseases
[…]